OpenStudy (anonymous):

Here's a fun problem: Integrate \[\int_{-\pi}^\pi \ln(2+\cos x)\,dx\]

hartnn (hartnn):

2 \(\int_0^\pi ..\)

OpenStudy (anonymous):

a fun problem

OpenStudy (anonymous):

ans is 0

OpenStudy (anonymous):

I think

hartnn (hartnn):

\(= 2 \int_0^\pi \ln (2-\cos x) dx\)

OpenStudy (anonymous):

Not zero, this here is an even function.

OpenStudy (anonymous):

\[\int\limits_{-a}^{a}f(x)dx=\int\limits_{0}^{2a}f(2a-x)dx\]

OpenStudy (anonymous):

did I mess up??

ganeshie8 (ganeshie8):

im thinking of parameter \[I(a) = \int_{-\pi}^\pi \ln(2+\cos(a x))\,dx\]

OpenStudy (anonymous):

hey @ganeshie8

OpenStudy (anonymous):

SITH WHATS THE ANSWR???

OpenStudy (anonymous):

@ganeshie8 I've been making slow progress with a slightly different parameter, \[I(b)=\int_{-\pi}^\pi \ln(2+b\cos x)\,dx\] There's definite an answer though, I'm curious to see its closed form.

OpenStudy (anonymous):

*if it has one

OpenStudy (anonymous):

DUDE I got it

OpenStudy (anonymous):

Here what Mathematica gave as an answer \[ -2 \pi \ln \left(4-2 \sqrt{3}\right) \]

OpenStudy (anonymous):

@eliassaab Mma won't always be there to help ;)

OpenStudy (anonymous):

FORGET MY PREVIOUS COMMENT

OpenStudy (anonymous):

I know, but just to check your answer if you get one

OpenStudy (anonymous):

but I am comin close

OpenStudy (anonymous):

For those interested, my parameterization gives me the following integral along the way: \[\int\frac{1}{b}\sqrt{\frac{2+b}{2-b}}(2+b)\,db\] which I'm told (thanks to WA) that there is an elementary antiderivative.

OpenStudy (anonymous):

conjugate

OpenStudy (anonymous):

\[\int \frac{\sqrt{\frac{b+2}{2-b}} (b+2)}{b} \, db=\\ \frac{\sqrt{\frac{b+2}{2-b}} \left(\sqrt{b+2} (b-2)+8 \sqrt{2-b} \sin ^{-1}\left(\frac{\sqrt{b+2}}{2}\right)-4 \sqrt{2-b} \tanh ^{-1}\left(\frac{\sqrt{b+2}}{\sqrt{2-b}}\right)\right)}{\sqrt{b+2}} \]

OpenStudy (anonymous):

whoa

OpenStudy (anonymous):

hey ARE U SURE ITS NOT LOG(1-cosx)???

OpenStudy (anonymous):

or log(1+cosx)??

OpenStudy (anonymous):

cmon man pls

OpenStudy (anonymous):

I solved this one

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

Yes, I came up with the problem. \(\ln(1\pm\cos x)\) is a bit strange, WA says both integrals converge despite the singularities: \(\ln(1+\cos x)\): http://www.wolframalpha.com/input/?i=Integrate%5BLog%5B1%2BCos%5Bx%5D%5D%2C%7Bx%2C-pi%2Cpi%7D%5D \(\ln(1-\cos x)\): http://www.wolframalpha.com/input/?i=Integrate%5BLog%5B1-Cos%5Bx%5D%5D%2C%7Bx%2C-pi%2Cpi%7D%5D

OpenStudy (anonymous):

I SOLVED IT

OpenStudy (anonymous):

U GET \[2\pi{\log1/2}\]

OpenStudy (anonymous):

I dont use wolfram lmao

OpenStudy (anonymous):

YA BCOZ when u use 2 .,.............u get stuck at 4-\[\cos^2x\]

OpenStudy (anonymous):

SITH is ur problem solvable???...what does wolfram say??

OpenStudy (anonymous):

Your answer is incorrect, the area of the region described by the integral lies above the x-axis, it can't possibly be negative \(\bigg(\)or zero, depending on if you meant \(2\pi\log\dfrac{1}{2}\) or \(\left.2\pi\dfrac{\log1}{2}\right)\).

OpenStudy (anonymous):

NOPE ITS NOT................IF U USE 1 instead of 2

OpenStudy (anonymous):

wait lemme check for errors

OpenStudy (anonymous):

The problem isn't integrating \(\ln(1+\cos x)\)...

OpenStudy (anonymous):

SITH..........I QUIT FROM THIS FUN PROBLEM.........

OpenStudy (anonymous):

Interestingly,\[\int_{-\pi}^\pi\ln(1+\cos x)\,dx=\int_{-\pi}^\pi\ln(1-\cos x)\,dx\] but that's irrelevant, seeing as there's a \(\bf2\) in the original problem.

OpenStudy (anonymous):

@ganeshie8 upon differentiating your parameterization I think we'll have to implement some complex contour, I remember seeing something of this form before: \[I'(a)=\int_{-\pi}^\pi\frac{-x\sin ax}{2+\cos ax}\,dx\]

ganeshie8 (ganeshie8):

yeah i got scared of that and trying if power series works smooth let me pause on power series a bit, this does look interesteing : \[\int_{-\pi}^\pi\ln(2+\cos x)\,dx=\int_{-\pi}^\pi\ln(2-\cos x)\,dx\] this is tryue because of that f(a-x) thingy.. messing wid this a bit we get \[I = \int_0^{\pi} \ln(4-\cos^2x) dx\] then its the dead end i guess

OpenStudy (anonymous):

Kudos to @eliassaab for the elegant solution, as seen here: http://openstudy.com/users/eliassaab#/updates/54a358eae4b0b8a54fb07d0b#/updates/54a358eae4b0b8a54fb07d0b